A pool ball leaves a 0.60-meter high table with an initial velocity of 2.4 m/s. Predict the time required for the pool ball to fall to the ground and the horizontal distance between the table's edge and the ball's landing location.


1
Expert's answer
2022-04-27T13:33:54-0400

The time required for the pool ball to fall to the ground

"t=\sqrt{2H/g}\\
=\sqrt{2*0.60/9.8}=0.35\:\rm s"

The horizontal distance between the table's edge and the ball's landing location

"L=v_0t=2.4*0.35=0.84\:\rm m"